The expression for the number of molecules in a Maxwellian gas having speeds in

the range v to v + dv is


dNV=4(3.14)N((M/(2(3.14)kBT))3/2 Vexp[-(mv2/2kBT)]dV)


Using this relation, obtain an expression for average speed. Also, plot Maxwellian

distribution function versus speed at three different temperatures.


1
Expert's answer
2021-03-31T07:18:39-0400

"f(v)dv=(\frac{m}{2\pi k T})^{\frac 32}4\pi v^2 e^{-\frac{mv^2}{2kT}}dv,"

"\frac{df(v)}{dv}=-8\pi(\frac{m}{2\pi k T})^{\frac 32}v e^{-\frac{mv^2}{2kT}}(\frac{mv^2}{2kT}-1)=0,"

"\frac{mv^2}{2kT}-1=0,"

"v=\sqrt{\frac{2kT}{m}}=\sqrt{\frac{2RT}{M}}."